OpenStudy (anonymous):

5(2x - 8) + 15 = -15 please explain steps

OpenStudy (anonymous):

first you do the distributive property >>> 10x - 40 + 15 = -15

OpenStudy (anonymous):

then you combine like terms >>>> 10x (-40 +15) = -15 >>>> 10x -25 = -15

OpenStudy (puppylover52):

well 2x-8=-16 i think

OpenStudy (anonymous):

and them combine like terms over the equal sign >>> 10x (-25+25) = (-15+25) >>>>> 10x = 10

OpenStudy (anonymous):

Then divide >>>> 10x / 10 = 10/10 >>>> x=1